118 4. In each case below, a solution is known to contain one of the two compounds listed. Suggest a single reagent, based on tests used in this experiment, to identify the compound. State the reagent to be added, the observation, and the conelusion a) KOH and KCI b) NH,Cl and KCI c) Fe(NOs)s and Al(NO,)s d) K CrOs and KNO 5. A solution is known to contain one of the following ions: Hg2,Sr, Cs . Refer to a solubility rules table from your text to suggest how the identity might be determined.
Answer #1

Question 4.

a)

KOH = strong base, will have a high pH, if we add HCl, expect neutralization, heat must be released, i.e. T will increase

KCl = this is ionic salt K+ and Cl- will be in solution, Add Hg2+ ions in order to precipitate HgCl2. If solid forms, then this is ionic.

Also, flame test shoudl show K+ ion color in flames as purple

b)

NH4Cl = this is slightly acidic, should react with any base to form NH3(g) + OH- + H+ = H2O

KCl = use flame test, if violet flame is seen, this contains K+ ions

c)

Fe(NO3)3

Fe3+ ions can be detected by:

Fe(NO3)3 + 6 KSCN = K3[Fe(SCN)6] + 3 KNO3

the solution complex will form a dark -red color

For Al(NO3)

Al3+ is present, it must be in basic media so:

The aluminum ion gets involved the adsorption of the red dye aluminon (which is actually the species - aurin tricarboxylic acid).

The Al(OH)3 will be seen as a precipitate with this dye.

d)

K2CrO4 = 2K+ + CrO4-2

Add Ag+ ions, they will preicpitate with

2Ag+ + CrO4-2 = Ag2CrO4(s)

if no precipitate forms, this should be the KNO3 solution.

Use flame test to identify its color (violet) and presence